OpenStudy (anonymous):

Find exact value of the following logarithm: Find exact value of the following logarithm: @Mathematics

OpenStudy (anonymous):

\[\log_{6}\sqrt[3]{6}\]

OpenStudy (id21):

\[\log_66^{\frac{1}{3}}=\frac{1}{3}\times 1=\frac{1}{3}\]